# Classification of connection graphs of global attractors for $S^{1}$-equivariant parabolic equations

### v2

@@ 3 Discussion and concluding remarks
As an application ot Theorem 1 we obtain the following result on connection graphs of reversible global attractors
Theorem 2.
Let $f=f(u,u_{x})$ and $g=g(u,u_{x})$ denote two reversible nonlinearities in the same lap signature class. Then the corresponding global attractors are connection equivalent,
(3.1) $\mathcal{A}_{f}\sim\mathcal{A}_{g}\ .$
Here (3.1) means connection equivalence between global attractors, [9, 10, 31]. This result is not immediate because all the critical elements in $\mathcal{P}$ fail the hyperbolicity condition (H) which would entail the transversality of stable and unstable manifolds and the strong Morse-Smale property. Hence, we replace stable and unstable manifolds by center stable and center unstable manifolds, and prove their automatic transverslity following closely the previous transversality results established for the general problem (1.1), see [11, 7]. See also [19, 1] for the original transversality results in the case of separated boundary conditions.
As already mentioned, every $2\pi$-periodic solution in $\mathcal{P}$ is a nonhomogeneous stationary wave frozen in time. So, let $\mathcal{F}$ denote the set of nonhomogeneous stationary waves, and $\mathcal{H}$ the set of heteroclinic orbits. Then, for the restrictive set of reversible nonlinearities $f\in\mathcal{R}$, the global attractor $\mathcal{A}_{f}$ of (1.1) decomposes as
(3.2) $\mathcal{A}_{f}=\mathcal{E}\cup\mathcal{F}\cup\mathcal{H}\ .$
By reversibility, each nonhomogeneous stationary solution $v(\cdot)\in\mathcal{F}$ generates a continuum of shifted copies $v(\cdot+\vartheta),\vartheta\in S^{1}$, all normally hyperbolic. This implies that each $v(\cdot)\in\mathcal{F}$ has a unique center manifold composed by its frozen shifted copies. In view of the gradient flow variational character of (1.1) for $f\in\mathcal{R}$, (see [8]), a non-stationary orbit $u(t,\cdot)$ converges either to equilibria or stationary waves as $t\rightarrow\pm\infty$. Hence, these orbit connections are essentially heteroclinic connections between equilibria, and the proof of transversality is similar to previous proofs not involving periodic orbits. Moreover, the proof of this transversality was already sketched in Propositions 3.1 and 3.2 of [11]. In the following, for each $v(\cdot)\in\mathcal{F}$, we let
Finally, we recall that a convenient notion of adjacency between two critical elements in $\mathcal{E}\cup\mathcal{F}$ establishes the existence of heteroclinic orbits, see Theorems 1.3 and 1.4 of [11].
We say that two (different) stationary waves $v_{\pm}(\cdot)\in\mathcal{F}$ are connected by a heteroclinic orbit $u(t,\cdot)$ if this connecting orbit converges to suitably phase shifted stationary waves $w_{\pm}:=v_{\pm}(\cdot+\vartheta_{\pm})$ as $t\rightarrow\pm\infty$.
Definition 1: We say that $W^{cu}(v_{-})$ and $W^{cs}(v_{+})$ are transverse,
@@
[…]
Let $u_{0}=u(0,\cdot)$ denote the initial condition of the connecting orbit. Then, if $S(t):X\rightarrow X$ denotes the semiflow generated by (1.1), we have that $u(t,x)=S(t)u_{0}(x)$. Moreover, if $DS(t)$ is the Fréchet derivative of $S(t)$, then for any $v_{0}\in X$ the curve $v(t,\cdot)=\bigl{(}DS(t)u_{0}(\cdot)\bigr{)}v_{0}(\cdot)$ in $X$ defines the classical solution $v(t,x)$ of the linearized equation
(3.5) $v_{t}=v_{xx}+f_{p}(u,u_{x})v_{x}+f_{u}(u,u_{x})v\ ,\quad x\in S^{1}\ ,\ t>0\ ,\quad v(0,x)=v_{0}(x)\ .$
Let $\{\lambda_{0}>\lambda_{1}>\lambda_{2}>\dots\}$ denote the eigenvalues of the linear second order differential operator in (3.5). If $v\in\mathcal{F}$ is a nonhomogeneous stationary wave, then by normal hyperbolicity we define its Morse index $i(v)$ as the number $k$ of strictly positive eigenvalues $\lambda_{j}$. We also have that $\lambda_{k+1}=0$ corresponds to the eigenfunction $v_{k+1}=v_{x}$. This shows that
Notice that, if we restrict $S^{1}$ to the Neumann interval $[0,\pi]$, then by Sturm-Liouville we conclude that the zero number of the eigenfunction $v_{j}$ is $z(v_{j})=2j$. Moreover, we recall that the zero number of $v(t,\cdot)$,
(3.7) $t\mapsto z(v(t,\cdot))\ ,$
is monotone nonincreasing (see Lemma 2.6 [25]), with strict decreasing at multiple zeros, (see [2]).
For the proof of transversality we follow almost verbatin the proof in [2]. We have the following zero number characterization of stable and unstable manifolds, adapted from Lemma 3 and Lemma 4 of [2]:
(i) Let $v\in\mathcal{F}$ have a positive Morse index $i(v)>0$. Then for each $v_{0}\in W^{u}(v)$ and $k=1,\dots,i(v)$ there exists a $k$-dimensional linear subspace $L_{k}(v_{0})$ of the tangent space of $W^{u}(v)$ at $v_{0}$, $T_{v_{0}}W^{u}(v)$, such that for each $w\in L_{k}(v_{0})\setminus\{0\}$ we have $z(w)<2k$. In particular, if $w\in T_{v_{0}}W^{u}(v)$ then $z(w)<2i(v)$.
(ii) Similarly, let $v_{0}\in W^{s}(v)$ and $w\in T_{v_{0}}W^{s}(v)\setminus\{0\}$. Then $z(w)\geq 2i(v)$.
Notice that this characterization immediately precludes the existence of homoclinic connecting orbits. So, we have the following automatic transversality result
Theorem 3.
Let $v_{\pm}\in\mathcal{F}$ denote normally hyperbolic fixed points of the semigroup $S(t)$ defined by (1.1). Then the center unstable manifold of $v_{-}$, $W^{cu}(v_{-})$, and the center stable manifold of $v_{+}$, $W^{cs}(v_{+})$, intersect transversely:
[…]
Proof of Theorem 3: Recall that if $W^{cu}(v_{-})\cap W^{cs}(v_{+})=\emptyset$ then, by definition, $W^{cu}(v_{-})$ and $W^{cs}(v_{+})$ are transverse. So, we assume an intersection point $u_{0}$ between the center unstable manifold of $v_{-}$ and the center stable manifold of $v_{+}$,
(3.9) $u_{0}\in W^{u}(w_{-})\cap W^{s}(w_{+})\subset W^{cu}(v_{-})\cap W^{cs}(v_{+})\ .$
Since the connecting orbit $u(t,\cdot)$ with $u(0,\cdot)=u_{0}$ lies in $W^{u}(w_{-})\cap W^{s}(w_{+})$, the invariance of these manifolds implies that
(3.10) $u_{t}(t,\cdot)\in T_{u_{0}}W^{u}(w_{-})\cap T_{u_{0}}W^{s}(w_{+})\ .$
Moreover, since $u_{t}(0,\cdot)$ is nonzero, by the zero number characterization of the stable and unstable manifolds we obtain
(3.11) $2i(v+)\leq z(u_{t}(0,\cdot)<2i(v_{-})\ .$
Let $N=i(v_{-})$. Then by the zero number characterization (i) there exists a $N$-dimensional linear subspace $L_{N}(u_{0})\subset T_{u_{0}}W^{u}(w_{-})$ such that for any nonzero $w\in L_{N}(u_{0})$ one has
(3.12) $z(w)<2N\ .$
On the other hand, the nonincreasing property of the zero number implies that for any nonzero $w\in T_{u_{0}}W^{s}(w_{+})$ we obtain
(3.13) $z(w)\geq 2N\ .$
This shows that
(3.14) $L_{N}(u_{0})\cap T_{u_{0}}W^{s}(w_{+})=\{0\}\ ,$
and since
(3.15) $\mathop{\mathrm{dim}}\nolimits L_{N}(u_{0})=N=\mathop{\mathrm{codim}}\nolimits T_{u_{0}}W^{s}(w_{+})\ ,$
we obtain
(3.16) $L_{N}(u_{0})\oplus T_{u_{0}}W^{s}(w_{+})=X\ .$
Hence we have
(3.17) $T_{u_{0}}W^{u}(w_{-})+T_{u_{0}}W^{s}(w_{+})=X$
[…]
Theorem 2 is a contribution to the geometric and qualitative theory of dynamical systems generated by parabolic partial differential equations. This result, applied to one-dimensional scalar semilinear PDEs, brings about the classification of these dynamical systems, already initiated in [11, 30, 12, 13]. Similar results are also available in the more complex case of monotone feedback delay differential equations. For results and references see [21, 22, 28].
@@
where $K$ again denotes the number of cyclicity regions, $k_{j}$ denotes the number of $2\pi$-periodic orbits in the $j$th annular or punctured disk region (this may be empty, in which case $k_{j}=0$), and $\ell_{i}^{j}$ denotes the period lap number of the $i$th $2\pi$-periodic orbit in the $j$th cyclicity region. The regular parenthesis structure represents the nesting of the periodic orbits. To illustrate this representation, we exhibit the lap signature class for the example of Fig. 6 (with $n=5$, $l=1$, $k_{1}=2$ and $k_{2}=k_{3}=1$) completing the parenthesis order structure shown in (2.15):
@@
The present result also shows that in the class of $S^{1}$-equivariant nonlinearities $f=f(u,u_{x})$ there is a continuous family of functions $f_{s},s\in[0,1]$, connecting the global Sturm attractors $\mathcal{A}_{f_{0}}=\{{\bf 0}\}$ and $\mathcal{A}_{f_{1}}$ through a finite number of local (degenerate) bifurcations. Here $s$ denotes the bifurcation parameter. Moreover, these bifurcations consist only of pitchfork and Hopf bifurcations. Notice that the bifurcations shown are degenerate because the family $f_{s}\in\mathcal{R},s\in[0,1]$, is entirely constructed in the symmetry class of reversible nonlinearities $f=f(u,u_{x})$, which implies its integrability. We also illustrate this bifurcation diagram in the example of Fig. 6. We obtain a family $f_{s},s\in[0,1]$, with six bifurcations points at $s=k/7,1\leq k\leq 6$, and seven (degenerate) structuraly stable regions $[0,1]\setminus\bigl{(}\cup_{1\leq k\leq 6}\{k/7\}\bigr{)}$. In Fig. 8 we illustrate the period maps in these seven regions. Since we are restricted to the integrable case $f_{s}\in\mathcal{R}$, all the bifurcations are degenerate. For general references see [5, 6, 15].
@@
In Fig. 9 we then show the bifurcation diagram from $\mathcal{A}_{0}=\{0\}$ to $\mathcal{A}_{1}$ and the connection graph of a nondegenerate global attractor $\widetilde{\mathcal{A}}_{1}$ of (1.1) homotopic to the (degenerate) $\mathcal{A}_{1}$. For this we extend the family $f_{s}$ to $s>1$, connecting the reversible (degenerate) $\mathcal{A}_{1}$ to a global attractor $\widetilde{\mathcal{A}}_{1}$ which is not reversible, see [12]. In this case $\widetilde{\mathcal{A}}_{1}$ is nondegenerate and we have the usual notions of stability of equilibria and rotation waves, [13]. See [11], Proposition 3.1, and [13], Theorem 10.
@@
Figure 9: Left: The bifurcation diagram for the reversible family $f_{s}\in\mathcal{R},s\in[0,1]$, (see Fig. 8). The bifurcation points are, respectively: a pitchfork of a stable saddle point (black ball); degenerate Hopf bifurcations (white balls); and, a pitchfork of centers (white square). Moreover, the sequence of bifurcations leading to $\mathcal{A}_{1}$ are: $(i)$ a supercritical pitchfork bifurcation of two stable saddle points at $s=1/7$; $(ii)$ a supercritical Hopf bifurcation of an unstable $2\pi$-periodic orbit at $s=2/7$; $(iii)$ a supercritical Hopf bifurcation of an unstable $2\pi$-periodic orbit at $s=3/7$; $(iv)$ a supercritical pitchfork bifurcation of two centers at $s=4/7$: $(v)$ a supercritical Hopf bifurcation of an unstable $2\pi$-periodic orbit at $s=5/7$ in the upper branch center; and, $(v)$ a supercritical Hopf bifurcation of an unstable $2\pi$-periodic orbit at $s=6/7$ in the lower branch center. At the right margin of the bifurcation diagram we indicate the Morse indices corresponding to the equilibria and rotating waves of the nondegenerate Neumann section of the global attractor $\widetilde{\mathcal{A}}_{1}$. Right: The connection graph of the nondegenerate global attractor $\widetilde{\mathcal{A}}_{1}$.
@@
The global attractor $\widetilde{\mathcal{A}}_{1}$ has three stable equilibria, denoted by $\{{\bf e}_{1},{\bf e}_{3}={\bf 0},{\bf e}_{5}\}$, and two unstable equilibria, $\{{\bf e}_{2},{\bf e}_{4}\}$, where
@@
(3.22) $\mathop{\mathrm{dim}}\nolimits\widetilde{\mathcal{A}}_{1}=\max\Bigl{\{}\mathop{\mathrm{dim}}\nolimits W^{u}({\bf e}_{j}),1+\mathop{\mathrm{dim}}\nolimits W^{u}({\bf v}_{k}),1\leq j\leq 5,1\leq k\leq 4\Bigr{\}}=3\ ,$
the global attractor $\widetilde{\mathcal{A}}_{1}$ is three dimensional. The heteroclinic orbit connections follow from the connections on the Neumann section of $\widetilde{\mathcal{A}}_{1}$, [11, 13]. The rotating waves, each one rotating with its own fixed speed $c_{j}$, appear on the slow stable/unstable invariant manifolds of the equilibria $\{{\bf e}_{2},{\bf e}_{3},{\bf e}_{4}\}$.
@@
where $\lambda\neq\lambda_{k}:=k^{2},k\in\mathbb{N}$, and $\varepsilon>0$. The set of equilibria of (3.23), =$\mathcal{E}=\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{3}\}$, has the equilibria ${\bf e}_{1}=+\sqrt{\lambda}$ and ${\bf e}_{3}=-\sqrt{\lambda}$, which are defined and hyperbolic for all $\lambda>0$. These are saddle points of (1.2). The third equilibrium, ${\bf e}_{2}=0$, is also hyperbolic for $\lambda\neq 0,\lambda_{k}$ and is a center of (1.2).
@@
Figure 10: Connection graph of the global attractor $\mathcal{A}_{\lambda}$ of (3.23) for $k^{2}<\lambda<(k+1)^{2}$. The numbers close to the edges are the corresponding Morse indices. This shows that $\mathop{\mathrm{dim}}\nolimits\mathcal{A}_{\lambda}=2k+1$.
Acknowledgments: The author acknowledges the continuous support of Bernold Fiedler (Freie U. Berlin). I am deeply thankful for his longtime friendship and collaboration. I am also gratefull to the continuous support of Isabel Rocha. Without her constant support, this endeavor would not be possible. This work was partially supported by FCT/Portugal through the projects UID/MAT/04459/2019 and UIDB/04459/2020.

### v5

@@ 4 Discussion and concluding remarks
Theorem 1 is a contribution to the geometric and qualitative theory of dynamical systems generated by parabolic partial differential equations. This result, applied to one-dimensional scalar semilinear PDEs, extends to the case of reversible nonlinearities the classification of these dynamical systems, already initiated in [14, 34, 15, 16]. Similar results are also available in the more complex case of monotone feedback delay differential equations. For results and references see [25, 26, 32].
@@
where $K$ again denotes the number of cyclicity regions, $k_{j}$ denotes the number of $2\pi$-periodic orbits in the $j^{th}$ annular or punctured disk region (this may be empty, in which case $k_{j}=0$), and $\ell_{i}^{j}$ denotes the period lap number of the $i^{th}$ $2\pi$-periodic orbit in the $j^{th}$ cyclicity region. The regular parenthesis structure represents the nesting of the periodic orbits. To illustrate this representation, we exhibit the lap signature class for the example of Fig. 6 (with $n=5$, $r=1$, $k_{1}=2$ and $k_{2}=k_{3}=1$) completing the parenthesis order structure shown in (3.15):
@@
The present result also shows that in the class of $S^{1}$-equivariant nonlinearities $f^{s}(u,u_{x})=f(s,u,u_{x}),s\in[0,1]$ there is a continuous family of functions $f^{s},s\in[0,1]$, connecting the global Sturm attractors $\mathcal{A}_{f^{0}}=\{{\bf 0}\}$ and $\mathcal{A}_{f^{1}}$ through a finite number of local (degenerate) bifurcations. Here $s$ denotes the bifurcation parameter. Moreover, these bifurcations consist only of pitchfork and Hopf bifurcations. Notice that the bifurcations shown are degenerate because the family $f^{s}\in\mathcal{R},s\in[0,1]$, is entirely constructed in the symmetry class of reversible nonlinearities $f=f(u,u_{x})$, which implies its integrability. We also illustrate this bifurcation diagram in the example of Fig. 6. We obtain a family $f^{s},s\in[0,1]$, with six bifurcations points at $s=k/7,1\leq k\leq 6$, and seven (degenerate) structuraly stable regions $[0,1]\setminus\bigl{(}\cup_{1\leq k\leq 6}\{k/7\}\bigr{)}$. In Fig. 8 we illustrate the period maps in these seven regions. Since we are restricted to the integrable case $f^{s}\in\mathcal{R}$, all the bifurcations are degenerate. For general references see [6, 8, 18].
@@
In Fig. 9 we then show the bifurcation diagram from $\mathcal{A}_{0}=\{0\}$ to $\mathcal{A}_{1}$ and the connection graph of a nondegenerate global attractor $\mathcal{A}_{s>1}$ of (1.1) homotopic to the (degenerate) $\mathcal{A}_{1}$. For this we extend the family $f^{s}$ to $s>1$, connecting the reversible (degenerate) $\mathcal{A}_{1}$ to a global attractor $\mathcal{A}_{s>1}$ which is not reversible, see [15]. In this case, $\mathcal{A}_{s>1}$ is nondegenerate and we have the usual notions of stability of equilibria and rotation waves, [16]. See [14], Proposition 3.1, and [16], Theorem 10.
@@
Figure 9: Left: The bifurcation diagram for the reversible family $f^{s}\in\mathcal{R},s\in[0,1]$, (see Fig. 8). The bifurcation points are, respectively: a pitchfork of a stable saddle point (black ball); degenerate Hopf bifurcations (white balls); and, a pitchfork of centers (white square). Moreover, the sequence of bifurcations leading to $\mathcal{A}_{1}$ are: $(i)$ a supercritical pitchfork bifurcation of two stable saddle points at $s=1/7$; $(ii)$ a supercritical Hopf bifurcation of an unstable $2\pi$-periodic orbit at $s=2/7$; $(iii)$ a supercritical Hopf bifurcation of an unstable $2\pi$-periodic orbit at $s=3/7$; $(iv)$ a supercritical pitchfork bifurcation of two centers at $s=4/7$: $(v)$ a supercritical Hopf bifurcation of an unstable $2\pi$-periodic orbit at $s=5/7$ in the upper branch center; and, $(v)$ a supercritical Hopf bifurcation of an unstable $2\pi$-periodic orbit at $s=6/7$ in the lower branch center. At the right margin of the bifurcation diagram we indicate the Morse indices corresponding to the equilibria and rotating waves of the nondegenerate Neumann section of the global attractor $\mathcal{A}_{s>1}$. Right: The connection graph of the nondegenerate global attractor $\mathcal{A}_{s>1}$.
@@
The global attractor $\mathcal{A}_{s>1}$ has three stable equilibria, denoted by $\{{\bf e}_{1},{\bf e}_{3}={\bf 0},{\bf e}_{5}\}$, and two unstable equilibria, $\{{\bf e}_{2},{\bf e}_{4}\}$, where
@@
(4.5) $\mathop{\mathrm{dim}}\nolimits\mathcal{A}_{s>1}=\max\Bigl{\{}\mathop{\mathrm{dim}}\nolimits W^{u}({\bf e}_{j}),1+\mathop{\mathrm{dim}}\nolimits W^{u}({\bf v}_{k}),1\leq j\leq 5,1\leq k\leq 4\Bigr{\}}=3\ ,$
the global attractor $\mathcal{A}_{s>1}$ is three dimensional. The heteroclinic orbit connections follow from the connections on the Neumann section of $\mathcal{A}_{s>1}$, [14, 16]. The rotating waves, each one rotating with its own fixed speed $c_{j}$, appear on the slow stable/unstable invariant manifolds of the equilibria $\{{\bf e}_{2},{\bf e}_{3},{\bf e}_{4}\}$.
@@
where $\lambda>0$ satisfies $\lambda\neq\lambda_{k}:=k^{2},k\in\mathbb{N}$, and $\varepsilon\neq 0$. The set of equilibria of (4.6), =$\mathcal{E}=\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{3}\}$, has the equilibria ${\bf e}_{1}=+\sqrt{\lambda}$ and ${\bf e}_{3}=-\sqrt{\lambda}$, which are defined and hyperbolic for all $\lambda>0$. These are saddle points of (1.2). The third equilibrium, ${\bf e}_{2}=0$, is also hyperbolic for $\lambda\neq\lambda_{1},\lambda_{2},\dots$, and is a center of (1.2).
@@
Figure 10: Connection graph of the global attractor $\mathcal{A}_{\lambda}$ of (4.6) for $k^{2}<\lambda<(k+1)^{2}$. The numbers close to the vertices are the corresponding Morse indices. This shows that $\mathop{\mathrm{dim}}\nolimits\mathcal{A}_{\lambda}=2k+1$.
As a concluding remark we point out that the nonlinearity $f=f(u,u_{x})$ can be regarded as a feedback control parameter for the reaction-diffusion equation (1.1). This provides a very interesting model for certain applications. We mention, for example, the possibility of stabilizing certain equilibria corresponding to reaction-diffusion patterns in chemical reactors with a proper choice of a nonlinearity $f=f(u,u_{x})$.
Acknowledgments: The author acknowledges the continuous support of Bernold Fiedler (Freie U. Berlin). I am deeply thankful for his longtime friendship and collaboration. I am also gratefull to the continuous support of Isabel Rocha. Without her constant support, this endeavor would not be possible. This work was partially supported by FCT/Portugal through the project UIDB/04459/2020 with DOI identifier 10-54499/UIDP/04459/2020.
The author also acknowledges a helpful discussion with Luis Barreira, and the expert comments of the three anonymous referees which largely contributed to the improvement of the manuscript.